# Aarti Ghatkesar

### v1

@@ Abstract
Achieving deep alignment between vision and language remains a central challenge for Multimodal Large Language Models (MLLMs). These models often fail to fully leverage visual input, defaulting to strong language priors. Our approach first provides insights into how MLLMs internally build visual understanding of image regions and then introduces techniques to amplify this capability. Specifically, we explore techniques designed both to deepen the model’s understanding of visual content and to ensure that these visual insights actively guide language generation. We demonstrate the superior multimodal understanding of our resultant model through a detailed upstream analysis quantifying its ability to predict visually-dependent tokens as well as $>$10 pt boost on visually challenging tasks.
@@ 3.4 Targeted Synthetic Data Generation
Figure 3: Examples of synthetic grid data for spatial awareness. Objects are randomly sampled from a large public image collection [4, 6, 7], and placed onto distinct locations within an N x N grid background.
@@ 4.2 Evaluation on SpatialEval Benchmark
We evaluate our proposed techniques on the SpatialEval [19] benchmark, designed to probe spatial understanding capabilities. Our results, summarized in Table 2, demonstrate a clear trend of performance improvement as our proposed components are incrementally introduced.
The addition of VisualLoss  strengthens the model’s fundamental understanding of visual content. While the direct gains from VisualLoss  are modest, its primary contribution is in enhancing the grounding necessary for subsequent components. This is evidenced when we introduce BlankTokens . We observe that the improved visual understanding provided by the visual loss amplifies the impact of the blank tokens, leading to substantial gains, including improvements of up to 5.5 percentage point on SpatialMap.
Adding synthetic data improves average performance across SpatialEvals tasks by over 9 percentage points compared to the baseline, despite minimal overlap with the benchmark tasks. This underscores the synthetic data’s effectiveness in enhancing fundamental visual grounding by leveraging our proposed training enhancements (Section 3).
Table 2: Impact of proposed techniques on SpatialEval accuracy - we observe consistent improvement as we layer in our innovations.
Technique Grid MazeNav Map Real
Baseline 31.0 27.5 27.1 48.9
VisualLoss
30.4 28.4 28.6 49.8
BlankTokens
31.9 30.6 32.6 50.7
+ Synthetic 41.1 29.5 51.4 51.1
VisualLossAdv
46.3 24.3 58.9 60.8
@@ 4.3 Future Extensions
This section presents preliminary explorations of extensions to our approach, inspired by recent community innovations, and highlight promising directions for future research.
Enhancing LLM backbone’s visual representation. We explore two extensions to further improve LLM’s internal visual representation. First, to mitigate the potential representational conflict arising from tasking the same LLM weights with processing both visual and textual concepts, we experimented with providing each modality with an independent set of weights [20, 8]. Cross-modal interaction was maintained through the standard self-attention mechanism. Our initial results in Table 1 indicate that this separation yields significantly enhances visual representation.
Next, we demonstrate that our approach can leverage advancements in image encoder to enables the LLM backbone to build richer visual representations. In particular, replacing I-JEPA [2] with the AM-Radio [12] encoder as the auxiliary image encoder for our visual loss (shown as VisualLossAdv  in Table 2) yields an additional gain of approximately 9.7 percentage points on the spatialreal subset (over 4 points on average across SpatialEval subsets).
These results provide strong evidence that our proposed training enhancements allow MLLMs to better leverage advancements in image encoders to build a richer internal visual representation. Further investigation of these extensions constitute key directions for our subsequent research.

### v2

@@ Abstract
Achieving deep alignment between vision and language remains a central challenge for Multimodal Large Language Models (MLLMs). These models often fail to fully leverage visual input, defaulting to strong language priors. Our approach first provides insights into how MLLMs internally build visual understanding of image regions and then introduces techniques to amplify this capability. We explore techniques designed both to deepen the model’s understanding of visual content and to ensure that these visual insights actively guide language generation. We demonstrate the superior multimodal understanding of our resultant model through a detailed upstream analysis on predicting visually-dependent tokens as well as $>$10 percentage point boost on a visually challenging task and a consistent boost across multiple tasks.
@@ 3.4 Targeted Synthetic Data Generation
Figure 3: Examples of synthetic grid data for spatial awareness. Objects are randomly sampled from a large public image collection [4, 6, 7], and placed onto distinct locations.
@@ 4.2 Evaluation on SpatialEval Benchmark
We evaluate our proposed techniques on the SpatialEval [19] benchmark111we had to enhance benchmark’s prompting and parsing capabilities which we open source as well for reproducibility, designed to probe spatial understanding capabilities. Our results, summarized in Table 2, demonstrate a clear trend of performance improvement across all benchmark subsets with our proposal.
The addition of VisualLoss  results in a significant improvement in model performance for SpatialMap and SpatialReal subsets of the benchmark. We attribute these improvements to VisualLoss ’s ability to strengthen the model’s understanding of visual content. Adding synthetic data provides consistent improvement across all benchmark subsets, despite minimal overlap with the benchmark tasks. In particular, it helps us narrow the performance gap with the Llama 3.2 11B model [5], even though our model is smaller (8B vs 11B), utilizes a lower input resolution (e.g., 224x224 vs Llama’s higher resolution of 560x560), and employs a simpler modality combination approach (input-layer feature appending vs multi-layer cross-attention). This underscores the synthetic data’s effectiveness in enhancing visual grounding by leveraging our proposed training enhancements (Section 3).
Table 2: Our proposed techniques consistently improve SpatialEvals accuracy across all benchmark subsets. Avg. is calculated as mean of normalized accuracy improvement.
Technique Grid MazeNav Map Real Avg. Baseline 34.5 15.1 44.1 50.4 1 + VisualLoss 31.3 17.0 50.7 56.3 1.075 + BlankTokens 30.1 22.9 47.1 52.6 1.125 + Synthetic 41.1 26.3 55.2 53.3 1.31 + VisualLossAdv 43.5 26.7 64.1 53.3 1.385 Llama 3.2 11B [5] 50.1 28.1 48.2 53.3 1.365
@@ 4.3 Future Extensions
We explore two extensions to further improve LLM’s internal visual representation. First, to mitigate the potential representational conflict arising from tasking the same LLM weights with processing both visual and textual concepts, we experimented with providing each modality with an independent set of weights [20, 8]. Cross-modal interaction was maintained through the standard self-attention mechanism. Our initial results in Table 1 (+ind. weights row) indicate that this separation significantly enhances visual representation - providing the biggest reduction in upstream loss thus far ($3.48$Rightarrow $3.20$).
Next, we demonstrate that VisualLoss  can leverage advancements in image encoders to help the LLM backbone build richer visual representations. We replace I-JEPA [2] with the AM-Radio [12] encoder as the auxiliary image encoder for our visual loss (shown as VisualLossAdv  in Table 2) yields an additional gain of approximately 9.1 percentage points on the SpatialMap subset. This enhancement enables our model to match the performance of Llama 3.2 11B on this task, despite our model being smaller, employing a simpler architecture, and using a lower-resolution visual input. Further investigation of these extensions constitute key directions for our subsequent research.